# From Fomo3D to Lottery DAPP: Analysis of Ethereum-Based Gambling Applications

### v1

@@ 3.1. Blockchain
Blockchain is a decentralized, distributed database composed of multiple interconnected blocks linked through cryptographic algorithms. Each block contains the hash value of the preceding block, transaction data, timestamps, and other relevant information (li2017discovering, ; kong2024characterizing, ). By leveraging consensus algorithms, the blockchain network achieves a consensus mechanism, ensuring data synchronization and eliminating the possibility of data forgery by any single node, thereby enabling a trustless system. Through mutually agreed protocols and smart contracts, nodes interact and compete autonomously, ensuring the system operates independently without human intervention. Cryptographic algorithms enable any participant to query data records via public interfaces while preventing data modification or repudiation (li2021hybrid, ). The chained structure facilitates efficient and rapid retrieval of transaction data, ensuring the traceability of data and transactions. Ethereum (john2025economics, ) is a decentralized open-source public blockchain platform with smart contract functionality. It features an integrated Turing-complete programming language, enabling users to develop DApps according to their specific requirements. As an application runtime platform, Ethereum ensures data transparency by making all data publicly accessible to nodes and immutable to third-party modifications. The Ethereum Virtual Machine (EVM) facilitates the execution and invocation of smart contracts. Additionally, Ethereum employs an account model, which reduces the cost of batch transaction processing, simplifies programming and development, and broadens the scope of application scenarios.
@@ 3.3. Consensus Mechanism
  • Proof of Work(PoW). PoW is the consensus mechanism initially used by Bitcoin and currently the most widely applied. Nodes, competing to solve a complex mathematical puzzle requiring computational power, gain the right and reward to generate new blocks upon success.
  • Proof of Stake(PoS). PoS is a consensus mechanism where a node’s influence in transaction validation and block generation is determined by the amount of digital assets it holds. To participate, nodes must stake a certain amount of their holdings.
@@
  • Delegated Proof of Stake(DPoS). DPoS is a consensus mechanism where asset holding stakeholders elect delegates to validate transactions and generate blocks. These delegates stake a certain amount of digital assets to secure the rights and rewards for block generation.
  • Practical Byzantine Fault Tolerance(PBFT). PBFT is a fault tolerant consensus mechanism. It addresses communication issues between nodes and prevents malicious attacks in distributed systems.
@@ 3.5. Dapp
DApp (izaguirre2024decentralized, ; kong2025uechecker, ) represents the integration of traditional applications (APPs) with blockchain technology, typically operating on a peer-to-peer (P2P) network. It serves as an enhancement and extension of conventional applications, with the key distinction being its decentralized nature. In DApp, participant information is either anonymous or protected, and operations are conducted through nodes on a peer-to-peer network. Smart contracts provide the foundational framework for decentralization, enabling trustless interactions between participants. From a practical perspective, a DApp can be succinctly conceptualized as a combination of smart contracts and traditional applications(li2025scalm, ), where smart contracts establish the prerequisites for decentralization. Structurally, DApps involve interactions between a front-end interface and users, as well as between smart contracts and the blockchain. This makes DApp publicly accessible programs that operate transparently on a network, leveraging blockchain’s inherent properties of immutability and traceability.
@@ 4.1. Fomo3D Game
Fomo3D (chen2019lottery, ), a blockchain-based cryptocurrency game on Ethereum, is a DApp named after the fear of missing out. Players buy FOMO tokens to join, adding funds to a communal pool. If no new purchase occurs in 24 hours, the last buyer wins the pool, encouraging rapid participation for higher profits. Fomo3D features an added $LastPlayerStanding$ mode, where each FOMO token purchase grants a keykey The last player holding a key wins the pool if no new purchases occur within 24 hours, heightening the game’s difficulty. The airdrop mechanism in Fomo3D rewards players who buy FOMO tokens with extra tokens or rewards, usually provided by the developers or sponsors, to boost participation. It has two types: regular airdrops, distributed at set times during the game to eligible players, and random airdrops, given unexpectedly to randomly selected players who have bought FOMO tokens. Airdrop terms and rewards are announced before the game starts. Generally, players need a minimum FOMO token holding to qualify, and rewards may change over time, so players should stay updated on game news and airdrop rules. A vulnerability in the game emerged in the airdrop mechanism.
@@ 4.3.1. Preventing the Prediction of ”Random Numbers” in Smart Contracts
(1) Use Real Random Number Generators. Traditional pseudo-random number generators are easy to crack by attackers. So, it’s advised to use real random number generators in smart contracts. These can get randomness from physical randomness (like quantum phenomena or thermal noise) or random events (such as mining or network data on the blockchain).
(2) Use Multiple Sources. Using multiple random number sources boosts security, as it’s tough for attackers to compromise them all. For instance, collect random numbers from different blockchain nodes or other data sources, and combine these sources to generate the final random number.
@@ 5.1.1. Traditional Server Client Model of APPs
The server client model is a special paradigm for web applications. It offers a static server for web apps, allowing clients to make dynamic requests to the server. In this model, clients only need to request services from the server. Upon receiving a request, the server creates a thread pool for the relevant task and sends the execution results back to the client. This approach ensures the server isn’t overloaded by multithreading operations. As shown in Figure 1.
@@ 5.2. Design and Implementation of Smart Contracts
Smart contracts on Ethereum are code snippets executed by the Ethereum Virtual Machine (EVM). They’re stored on the blockchain in Ethereum’s unique binary format, known as EVM bytecode. The EVM simplifies smart contract development compared to other blockchain systems. These contracts can be written in high level languages, and Ethereum currently supports three: Solidity, Serpent, and LLL.
@@ 6.1.1. MetaMask Wallet and Account
MetaMask is a popular browser extension for interacting with Ethereum DApps. It serves as a digital wallet for storing Ethereum and ERC-20 tokens. It also offers a secure interface for users to sign transactions, transfer tokens, and access DeFi protocols. A key feature is its ability to connect to various Ethereum networks, including the mainnet, testnets, and custom networks.
@@ 6.3. Deployment of Smart Contracts
The deployment of smart contracts involves publishing the contract code onto the blockchain so it can be accessed and executed within the blockchain network. This paper demonstrates deploying a smart contract to the Sepolia test network of Ethereum using a programming language like JavaScript. The process involves connecting to the Ethereum network with the Web3.js library and creating a smart contract object using the contract’s ABI and bytecode. The contract is then deployed using the deploydeploy method of the smart contract object, and the contract address is obtained upon successful deployment. Notably, the web3 instance should use the user’s provider. If the MetaMask plugin is installed in the browser, it automatically creates a web3 instance with the user’s provider. The provider can be accessed via window.web3.currentProvider and set in the application’s web3 instance. Key code snippets for this process are provided as Listing 7.

### v2

@@ 3.1. Blockchain
Blockchain is a decentralized, distributed database composed of multiple interconnected blocks linked through cryptographic algorithms. Each block contains the hash value of the preceding block, transaction data, timestamps, and other relevant information (li2017discovering, ; kong2024characterizing, ; wu2024blockchain, ). By leveraging consensus algorithms, the blockchain network achieves a consensus mechanism, ensuring data synchronization and eliminating the possibility of data forgery by any single node, thereby enabling a trustless system. Through mutually agreed protocols and smart contracts, nodes interact and compete autonomously, ensuring the system operates independently without human intervention. Cryptographic algorithms enable any participant to query data records via public interfaces while preventing data modification or repudiation (li2021hybrid, ). The chained structure facilitates efficient and rapid retrieval of transaction data, ensuring the traceability of data and transactions. Ethereum (john2025economics, ) is a decentralized open-source public blockchain platform with smart contract functionality. It features an integrated Turing-complete programming language, enabling users to develop DApps according to their specific requirements. As an application runtime platform, Ethereum ensures data transparency by making all data publicly accessible to nodes and immutable to third-party modifications. The Ethereum Virtual Machine (EVM) facilitates the execution and invocation of smart contracts. Additionally, Ethereum employs an account model, which reduces the cost of batch transaction processing, simplifies programming and development, and broadens the scope of application scenarios.
@@ 3.3. Consensus Mechanism
  • Proof of Work(PoW). PoW (maadallah2025enhancing, ; mahdi2024blockchain, ) is the consensus mechanism initially used by Bitcoin and currently the most widely applied. Nodes, competing to solve a complex mathematical puzzle requiring computational power, gain the right and reward to generate new blocks upon success.
  • Proof of Stake(PoS). PoS (john2025proof, ; pierro2024gas, ; li2024proof, ) is a consensus mechanism where a node’s influence in transaction validation and block generation is determined by the amount of digital assets it holds. To participate, nodes must stake a certain amount of their holdings.
@@
  • Delegated Proof of Stake(DPoS). DPoS (cao2024delegated, ; gupta2024profit, ; varadarajan2024innovative, ) is a consensus mechanism where asset holding stakeholders elect delegates to validate transactions and generate blocks. These delegates stake a certain amount of digital assets to secure the rights and rewards for block generation.
  • Practical Byzantine Fault Tolerance(PBFT). PBFT (li2024dynamic, ; zhou2024implementing, ) is a fault tolerant consensus mechanism. It addresses communication issues between nodes and prevents malicious attacks in distributed systems.
@@ 3.5. Dapp
DApp (izaguirre2024decentralized, ; kong2025uechecker, ; aufiero2024dapps, ) represents the integration of traditional applications (APPs) with blockchain technology, typically operating on a peer-to-peer (P2P) network (raza2024optimal, ; parhamfar2024towards, ) . It serves as an enhancement and extension of conventional applications, with the key distinction being its decentralized nature. In DApp, participant information is either anonymous or protected, and operations are conducted through nodes on a peer-to-peer network. Smart contracts provide the foundational framework for decentralization, enabling trustless interactions between participants. From a practical perspective, a DApp can be succinctly conceptualized as a combination of smart contracts and traditional applications(li2025scalm, ), where smart contracts establish the prerequisites for decentralization. Structurally, DApps involve interactions between a front-end interface and users, as well as between smart contracts and the blockchain. This makes DApp publicly accessible programs that operate transparently on a network, leveraging blockchain’s inherent properties of immutability and traceability.
@@ 4.1. Fomo3D Game
Fomo3D (chen2019lottery, ; yang2025fomo, ), a blockchain-based cryptocurrency game on Ethereum, is a DApp named after the fear of missing out. Players buy FOMO (zhang2025security, ) tokens to join, adding funds to a communal pool. If no new purchase occurs in 24 hours, the last buyer wins the pool, encouraging rapid participation for higher profits. Fomo3D features an added $LastPlayerStanding$ mode, where each FOMO token purchase grants a keykey The last player holding a key wins the pool if no new purchases occur within 24 hours, heightening the game’s difficulty. The airdrop mechanism in Fomo3D rewards players who buy FOMO tokens with extra tokens or rewards, usually provided by the developers or sponsors, to boost participation. It has two types: regular airdrops, distributed at set times during the game to eligible players, and random airdrops, given unexpectedly to randomly selected players who have bought FOMO tokens. Airdrop terms and rewards are announced before the game starts. Generally, players need a minimum FOMO token holding to qualify, and rewards may change over time, so players should stay updated on game news and airdrop rules. A vulnerability in the game emerged in the airdrop mechanism.
@@ 4.3.1. Preventing the Prediction of ”Random Numbers” in Smart Contracts
(1) Use Real Random Number Generators. Traditional pseudo-random number generators (naik2024review, ) are easy to crack by attackers. So, it’s advised to use real random number generators in smart contracts. These can get randomness from physical randomness (like quantum phenomena or thermal noise) or random events (such as mining or network data on the blockchain).
(2) Use Multiple Sources. Using multiple random number sources boosts security (alawida2024enhancing, ) , as it’s tough for attackers to compromise them all. For instance, collect random numbers from different blockchain nodes or other data sources, and combine these sources to generate the final random number.
@@ 5.1.1. Traditional Server Client Model of APPs
The server client model (nyabuto2024architectural, ) is a special paradigm for web applications. It offers a static server for web apps, allowing clients to make dynamic requests to the server. In this model, clients only need to request services from the server. Upon receiving a request, the server creates a thread pool for the relevant task and sends the execution results back to the client. This approach ensures the server isn’t overloaded by multithreading operations. As shown in Figure 1.
@@ 5.2. Design and Implementation of Smart Contracts
Smart contracts on Ethereum are code snippets executed by the Ethereum Virtual Machine (EVM) (bigiotti2024interoperability, ; song2024empirical, ; prifti2024smart, ) . They’re stored on the blockchain in Ethereum’s unique binary format, known as EVM bytecode. The EVM simplifies smart contract development compared to other blockchain systems. These contracts can be written in high level languages, and Ethereum currently supports three: Solidity, Serpent, and LLL.
@@ 6.1.1. MetaMask Wallet and Account
MetaMask (khan2024review, ) is a popular browser extension for interacting with Ethereum DApps. It serves as a digital wallet for storing Ethereum and ERC-20 tokens. It also offers a secure interface for users to sign transactions, transfer tokens, and access DeFi protocols. A key feature is its ability to connect to various Ethereum networks, including the mainnet, testnets, and custom networks.
@@ 6.3. Deployment of Smart Contracts
The deployment of smart contracts involves publishing the contract code onto the blockchain so it can be accessed and executed within the blockchain network. This paper demonstrates deploying a smart contract to the Sepolia test network of Ethereum using a programming language like JavaScript (rasti2025automated, ; valencia2025blockchain, ) . The process involves connecting to the Ethereum network with the Web3.js library and creating a smart contract object using the contract’s ABI and bytecode (antonino2024refinement, ) . The contract is then deployed using the deploydeploy method of the smart contract object, and the contract address is obtained upon successful deployment. Notably, the web3 instance should use the user’s provider. If the MetaMask plugin is installed in the browser, it automatically creates a web3 instance with the user’s provider. The provider can be accessed via window.web3.currentProvider and set in the application’s web3 instance. Key code snippets for this process are provided as Listing 7.